# 5- and 6-membered rings: A natural orbital functional study

### v1

@@ I Introduction
While the emergence of deep-learning and similar techniques has led to an improvement of parametrized methods, there is still room for ab-initio modern electronic structure methods. The latter are the unique alternative to practice discovery science, as recently shown by J. J. Eriksen et al.[1], who established the ground-state of Benzene by means of a blind challenge. However, emerging electronic structure methods require benchmarking, not only as a tool for comparison but a necessary test for validation. Benchmark studies provide a quantitative measure of the errors introduced by an approximation in computing different observables, which is essential for assessing the reliability of new approaches. Damour and co-workers [2] extended the aforementioned study of Benzene to a 12 molecular set compound by five- and six-membered rings. They investigated the performance and convergence properties of popular single-reference approaches, such as the Møller-Plesset perturbation series and the coupled-cluster (including iterative approximations) series, in comparison with full configuration interaction (FCI) correlation energy estimates. More importantly, the set included simple aromatic rings form the basis of more complex molecules of biological interest, so an accurate description is desired before going for larger and more complex systems. The motivation of the present study is to employ this molecular set to validate the performance of recent Natural Orbital Functional (NOF) approaches on molecules predominantly dynamic in correlation character.
@@
Despite these advances, NOFT remains underutilized, primarily due to two factors. First, NOF methods are not yet implemented in widely used electronic structure software packages. Although the open-source DoNOF program[34] for NOF calculations represents a significant step forward, broader integration is still needed. Second, accessible and systematic assessments of NOFs’ performance are scarce, making it difficult for researchers to gauge its reliability. In this vein, while the aforementioned GNOF approximation has been tested on strongly correlated models,[35, 36] its accuracy on systems dominated by dynamic correlation is undetermined yet, so a step forward in this direction is intended in the present work.
@@ II Electron-pairing-based NOFs
Recently, a modified version of GNOF, denoted GNOFm, reintroduces the interactions between strongly occupied orbitals in the antiparallel spin blocks, as originally proposed in PNOF7.[24, 25] This refinement has shown improved accuracy for describing the singlet triplet energy gaps along the linear n-acene series.[33] Within this framework, the inter-subspace static component takes the following compact form:
@@
The solution is established by optimizing the energy with respect to the ONs and NOs, separately. Therefore, orbitals vary along the optimization process until the most favorable orbital interactions are found. All calculations have been carried out using the DoNOF code [34] and the recently implemented orbital optimization algorithm.[33]
@@ III Motivation and Methodology
In Ref. [2], Damour et al. provided accurate FCI correlation energy estimates for twelve cases of five- and six-membered ring molecules, namely: Cyclopentadiene, Furan, Imidazole, Pyrrole, Thiophene, Benzene, Pyrazine, Pyridazine, Pyridine, Pyrimidine, s-Tetrazine, and s-Triazine. Hence, the set involves systems with atoms of the first to third lines of the periodic table. An schematic representation of the latter is shown in Fig. 1. In particular, Damour and co-workers reported optimized-orbital selected configuration interaction calculations for a correlation-consistent double-$\zeta$ Dunning basis set (cc-pVDZ),[55] as a reference for further studying the convergence of the Møller-Plesset perturbation theory series and the iterative approximate coupled-cluster series. Even in the context of the cc-pVDZ basis set, computing FCI result of these molecules is too computational demanding. Indeed, today carrying out coupled-cluster with singles, doubles, triples, and quadruples (CCSDTQ) calculations for molecules larger than benzene is prohibitively expensive or at least not practical.[1] This situation puts NOF approaches in an interesting position to run calculations employing larger basis sets from cc-pVDZ to cc-pV5Z. In the following, we use this molecular set to study GNOF and GNOFm correlation energies and their convergence with the size of the basis set. We provide complete-basis-set (CBS) estimates for these approximations, as well as for the ground-state gold standard coupled-cluster singles, doubles, and perturbative triples CCSD(T). The CCSD(T) calculations required a significant effort, especially in the largest cases at cc-pV5Z, which required around 1.4T of RAM, hence becoming infeasible for common computational configurations in contrast to NOF calculations. Following the work by Damour and co-workers, geometries of the molecular systems, obtained at the CC3/aug-cc-pVTZ level of theory, were extracted from Ref. [56].
@@
The DoNOF code [34] was employed for GNOF and GNOFm calculations, whereas CCSD(T) calculations were carried out with the PSI4 software package.[57] In contrast to Damour et al., no frozen core orbitals were considered in the present study. All electrons are correlated through all orbitals given in the basis set within the NOFT framework. The latter is, indeed, a strength of NOFs and their actual advantage with respect to typically used methods for multireference correlation, which require to define an active space where electrons are correlated. Finally, the resolution of identity approximation was used for integral evaluation in NOF calculations. [13] The latter was not employed in CCSD(T) calculations. However, as demonstrated[58] by DePrince III and Sherill, it would affect CCSD(T) energies, at most, in the order of a few $mE_{h}$, so in any case it alters neither the reported results nor the obtained conclusions.
@@ IV Results
In this section, we analyze GNOF and GNOFm correlation energies for the aforementioned set of molecules, and compare them with CCSD(T) calculations. Note that correlation energies refer to the difference between energies given by a correlated method $E_{\mathrm{M}}$ and the Hartree-Fock energies $E_{\mathrm{HF}}$, i.e. $E_{\mathrm{corr}}=E_{\mathrm{M}}-E_{\mathrm{HF}}$.
Correlation energies for GNOF, GNOFm, and CCSD(T) are shown in Fig. 2 for increasing size correlation-consistent Dunning basis sets (cc-pVXZ, X=2-5). Here, molecules are ordered from smaller to larger correlation energies, according to the numbering presented in Fig. 1. Note that for thiophene (no. 3) with the cc-pV5Z basis, the reported values are not the raw results obtained directly from each method; the cc-pV5Z calculations show an unwarranted drop not observed for the other molecules or for thiophene with the remaining Dunning basis sets (cc-pVXZ, X=2-4). We attribute this behavior to a limitation in the sulfur cc-pV5Z basis set design. Consequently, we report an interpolated estimate from the X = 2, 3, and 4 Dunning sets. A detailed discussion of this issue is given in the second section of the Supplementary Material.
A first look to the plot reveals that NOF and CCSD(T) curves are roughly parallel, represented by dashed and solid lines, respectively, so the molecular description agrees for both methods independently of the size of the basis set, as well as of the different studied molecular rings. Probably the most noticeable disagreement is obtained for benzene (no. 6) when the GNOF/cc-pVTZ methodology is employed (dotted red curve). Correlation energy corresponding to the latter converges too rapidly when using GNOF in comparison with GNOFm and CCSD(T), thus the energy difference between cc-pVDZ and cc-pVTZ is slightly larger for GNOF than for the latter. Interestingly, this difference is removed when GNOFm is employed, which reveals a similar result to that obtained with CCSD(T). GNOFm does not provide a too low correlation energy for Benzene by using the cc-pVTZ basis set, in contrast with GNOF, so the behaviour obtained for this molecule follows the line of other systems in the set. Overall, GNOFm retrieves more correlation energy than its predecessor GNOF, and less than reference CCSD(T) calculations. Nevertheless, in the case of small basis sets there are a few exceptions where GNOFm and CCSD(T) compare very accurately or the former provides larger correlation energies than the latter, as it can be seen in Fig. 2.
@@
Figure 2: Correlation energies ($E-E_{\mathrm{HF}}$) in mEh for the selected set of molecules, obtained by using GNOF (dotted lines), GNOFm (dashed lines), and CCSD(T) (solid lines) with the $c$-pVXZ basis sets, $X=2,3,4,5$ being the cardinal number of the basis set. Molecules ordered according to the numbering given in Fig. 1.
Figure 3: Correlation energies ($E-E_{\mathrm{HF}}$) in mEh for the selected set of molecules, obtained by using GNOF (dotted lines) and GNOFm (dashed lines) with the $c$-pVXZ basis sets ($X=2,3,4,5$), together with the resulting complete basis set (CBS) limit estimates. Details corresponding to the latter are given throughout the text, as well as in Fig. 4 and Table 1.
@@
1 Cyclopentadiene -1083.4 -1114.4 -1156.5
2 Pyrrole -1103.5 -1153.9 -1197.9
3 Thiophene -1140.3 -1173.9 -1228.6
4 Furan -1150.0 -1182.9 -1225.9
5 Imidazole -1152.6 -1188.5 -1236.8
6 Benzene -1229.5 -1302.7 -1359.7
7 Pyridine -1304.5 -1342.5 -1397.8
8 Pyrazine -1333.0 -1378.4 -1439.1
9 Pyrimidine -1334.8 -1386.7 -1435.2
10 Pyridazine -1332.9 -1380.8 -1443.4
11 Triazine -1363.7 -1416.0 -1470.5
12 Tetrazine -1414.3 -1465.1 -1528.6
@@
Table 1: Complete basis set (CBS) extrapolated correlation energies ($E-E_{\mathrm{HF}}$) in mEh for the 12 molecular systems, computed using GNOF, GNOFm, and CCSD(T). An exponential extrapolation scheme, $E(X)=E_{\mathrm{CBS}}+a_{1}\cdot\exp(-a_{2}X)$, was employed with $X=2,3,4,5$ as the cardinal number of the basis set. For Thiophene, the extrapolation was performed using $X=2,3,4$.
As shown in Fig. 3, energy differences for the same molecule decrease as the size of the basis sets is augmented. This suggests that we are approaching the CBS limit for the reported correlation energies. Therefore, we computed the CBS limit correlation energies for the twelve molecules set using GNOF and GNOFm. An exponential function like extrapolation scheme was employed to obtain the CBS limit, $E(X)=E_{\mathrm{CBS}}+a_{1}\cdot\exp(-a_{2}X)$, $X=2,3,4,5$ being the cardinal number of the basis set. Interestingly, the form of the curve barely changes from cc-pVQZ to cc-pV5Z and from cc-pV5Z to the CBS limit, with the exception of the Benzene molecule in the case of GNOF calculations, which has been discussed before. Therefore, NOF calculations rapidly converge with the increasing size of Dunning correlation-consistent basis sets, so including more orbitals in the calculations just means lowering the total energy for GNOF and GNOFm calculations beyond cc-pVQZ. Previous studies [59] demonstrated similar results for different extrapolation schemes within the NOFT framework. Convergence of CCSD(T) correlation energies with the increasing size of the basis set can be seen in Fig. S1 from the Supplementary Material. Qualitatively, there are no significant differences with respect to the convergence obtained for GNOF and GNOFm.
The corresponding GNOF and GNOFm CBS estimated values are given in Table 1, which also presents extrapolations CCSD(T), performed using the same procedure. An inspection of CBS limit molecular correlation energies reveals an agreement within 100 mEh for GNOF in most cases, values that are even improved to around 50 mEh when GNOFm is utilized. The results shown in Table 1 are summarized in Fig. 4. GNOFm energies systematically get closer to CCSD(T) results when the static term between electron pairs is modified according to Eq. (16). In other words, Fig. 4 reveals that GNOFm CBS correlation energies reduce differences between GNOF and CCSD(T) to the half. Additionally, as it is shown in Figs. 2 and 3, the improvement is obtained for all basis sets studied.
@@
Figure 4: Complete basis set (CBS) extrapolated correlation energies ($E-E_{\mathrm{HF}}$) in mEh for the 12 molecular systems, computed using GNOF, GNOFm, and CCSD(T). An exponential extrapolation scheme, $E(X)=E_{\mathrm{CBS}}+a_{1}\cdot\exp(-a_{2}X)$, was employed with $X=2,3,4,5$ as the cardinal number of the basis set. For Thiophene, the extrapolation was performed using $X=2,3,4$.
@@
Finally, Fig. 4 presents a direct comparison of CBS-extrapolated correlation energies (in mEh) for the 12 rings obtained with GNOF, GNOFm, and CCSD(T). Both GNOF and its recent modification GNOFm provide very accurate descriptions of the five- and six-membered rings, as their traces lie close to and largely parallel with the CCSD(T) reference. Notably, Fig. 4 also shows a quantitative improvement of GNOFm over GNOF, consistent with Lew-Yee and Piris.[33] More importantly, the figure demonstrates that NOFs recover dynamic-correlation effects across the entire family of correlation-consistent Dunning basis sets and for all systems considered, proving the robustness of the Global NOF approach. Unlike earlier approximations, these functionals incorporate dynamic correlation within the energy expression itself and therefore do not require perturbative corrections.

### v2

@@ I Introduction
While the emergence of deep-learning and similar techniques has led to an improvement of parametrized methods, there is still room for ab-initio modern electronic structure methods. The latter are the unique alternative to practice discovery science, as recently shown by J. J. Eriksen et al.[1], who established the ground-state of Benzene by means of a blind challenge. However, emerging electronic structure methods require benchmarking, not only as a tool for comparison but a necessary test for validation. Benchmark studies provide a quantitative measure of the errors introduced by an approximation in computing different observables, which is essential for assessing the reliability of new approaches. Benzene molecule has been lately exploited as a test system for new methods [2, 3, 4]. Damour and co-workers [5] extended the aforementioned study of Benzene to a 12 molecular set compound by five- and six-membered rings. They investigated the performance and convergence properties of popular single-reference approaches, such as the Møller-Plesset perturbation series and the coupled-cluster (including iterative approximations) series, in comparison with full configuration interaction (FCI) correlation energy estimates. More importantly, the set included simple aromatic rings form the basis of more complex molecules of biological interest, so an accurate description is desired before going for larger and more complex systems. The motivation of the present study is to employ this molecular set to validate the performance of recent Natural Orbital Functional (NOF) approaches on molecules predominantly dynamic in correlation character.
@@
Despite these advances, NOFT remains underutilized, primarily due to two factors. First, NOF methods are not yet implemented in widely used electronic structure software packages. Although the open-source DoNOF program[37, 38] for NOF calculations represents a significant step forward, broader integration is still needed. Second, accessible and systematic assessments of NOFs’ performance are scarce, making it difficult for researchers to gauge its reliability. In this vein, while the aforementioned GNOF approximation has been tested on strongly correlated models,[39, 40] its accuracy on systems dominated by dynamic correlation is undetermined yet, so a step forward in this direction is intended in the present work.
@@ II Electron-pairing-based NOFs
Recently, a modified version of GNOF, denoted GNOFm, reintroduces the interactions between strongly occupied orbitals in the antiparallel spin blocks, as originally proposed in PNOF7.[27, 28] This refinement has shown improved accuracy for describing the singlet triplet energy gaps along the linear n-acene series.[36] Within this framework, and assuming real orbitals so that $L_{pq}=K_{pq}$, the inter-subspace static component takes the following compact form:
@@
The solution is established by optimizing the energy with respect to the ONs and NOs, separately. Therefore, orbitals vary along the optimization process until the most favorable orbital interactions are found. All calculations have been carried out using the DoNOF code [37, 38] and the recently implemented orbital optimization algorithm.[36]
@@ III Motivation and Methodology
In Ref. [5], Damour et al. provided accurate FCI correlation energy estimates for twelve cases of five- and six-membered ring molecules, namely: Cyclopentadiene, Furan, Imidazole, Pyrrole, Thiophene, Benzene, Pyrazine, Pyridazine, Pyridine, Pyrimidine, s-Tetrazine, and s-Triazine. Hence, the set involves systems with atoms of the first to third lines of the periodic table. An schematic representation of the latter is shown in Fig. 1. In particular, Damour and co-workers reported optimized-orbital selected configuration interaction calculations for a correlation-consistent double-$\zeta$ Dunning basis set (cc-pVDZ),[59] as a reference for further studying the convergence of the Møller-Plesset perturbation theory series and the iterative approximate coupled-cluster series. Even in the context of the cc-pVDZ basis set, computing FCI result of these molecules is too computational demanding. Indeed, today carrying out coupled-cluster with singles, doubles, triples, and quadruples (CCSDTQ) calculations for molecules larger than benzene is prohibitively expensive or at least not practical.[1] This situation puts NOF approaches in an interesting position to run calculations employing larger basis sets from cc-pVDZ to cc-pV5Z. Nevertheless, as discussed below, all-electron calculations have been employed consistently throughout this work, with no frozen electrons. Therefore, basis sets including weighted core-valence functions are preferred. To this end, we have included Dunning’s weighted core-valence basis sets cc-pwCVXZ [60], whereas results obtained by using the cc-pVXZ basis set family are shown in the Supplementary Information. The latter are carried out for X=2-5 cardinal numbers, however, the X=5 level is omitted from the cc-pwCVXZ calculations due to the marked increase in the number of basis functions [61] and the computational cost. In the following, we use these molecular sets to study GNOF and GNOFm correlation energies and their convergence with the size of the basis set. We provide complete-basis-set (CBS) estimates for these approximations, as well as for the ground-state gold standard coupled-cluster singles, doubles, and perturbative triples CCSD(T). Following the work by Damour and co-workers, geometries of the molecular systems, obtained at the CC3/aug-cc-pVTZ level of theory, were extracted from Ref. [62].
@@
The DoNOF code [37, 38] was employed for GNOF and GNOFm calculations, whereas CCSD(T) calculations were carried out with the PSI4 software package.[63] In contrast to Damour et al., no frozen core orbitals were considered in the present study. All electrons are correlated through all orbitals given in the basis set within the NOFT framework. The latter is, indeed, a strength of NOFs and their actual advantage with respect to typically used methods for multireference correlation, which require to define an active space where electrons are correlated. Finally, the resolution of identity approximation was used for integral evaluation in NOF calculations. [16] The latter was also employed in CCSD(T) calculations through the density-fitting option available in the PSI4 code [63]. As demonstrated[64] by DePrince III and Sherill, this would affect CCSD(T) energies, at most, in the order of a few $mE_{h}$, so in any case it alters neither the reported results nor the obtained conclusions.
@@ IV Results
In this section, we analyze GNOF and GNOFm correlation energies for the aforementioned set of molecules, and compare them with CCSD(T) calculations. Note that correlation energies refer to the difference between energies given by a correlated method $E_{\mathrm{M}}$ and the Hartree-Fock energies $E_{\mathrm{HF}}$, i.e. $E_{\mathrm{corr}}=E_{\mathrm{M}}-E_{\mathrm{HF}}$. A summary of raw correlation energies is publicly available in Ref. [65].
Correlation energies for GNOF, GNOFm, and CCSD(T) are shown in Fig. 2 for increasing size weighted core-valence Dunning basis sets (cc-pwCVXZ, X=2-4). Here, molecules are ordered from smaller to larger correlation energies, according to the numbering presented in Fig. 1. In the case of thiophene, including $d$ functions in the basis sets is required to obtain a correct description due to the presence of a sulfur atom [66]. In terms of correlation energy, GNOF retrieves 1095 mEh and 1122 mEh for thiophene using the cc-pwCVDZ and aug-cc-pwCVDZ basis sets, respectively. However, this difference reduces to a few mEh for larger basis sets. Concretely, GNOF thiophene correlation energies read as 1397 mEh and 1398 mEh for cc-pwCVTZ and aug-cc-pwCVTZ, respectively, and 1500 mEh and 1505 mEh for cc-pwCVQZ and aug-cc-pwCVQZ, respectively.
A first look to Fig. 2 reveals that NOF and CCSD(T) curves are roughly parallel, represented by dashed and solid lines, respectively, so the molecular description agrees for both methods independently of the size of the basis set, as well as of the different studied molecular rings. In detail, results corresponding to the smallest basis set employed, cc-pwCVDZ, show that GNOF and GNOFm retrieve more correlation energy than CCSD(T), represented by the solid blue line. Nevertheless, this result is reversed when the basis set size is increased. In the case of cc-pwCVTZ and cc-pwCVQZ calculations, represented by red and green lines, respectively, NOF correlation energies remain above CCSD(T) results. GNOFm shows a better agreement with CCSD(T) than its predecessor GNOF in both cases. This agreement is particularly accurate when using the cc-pwCVTZ basis set. However, CCSD(T) correlation energies get larger than GNOFm ones when the basis set increases from cc-pwCVTZ to cc-pwCVQZ, and thereby GNOFm differences with respect to CCSD(T) grow with the basis set. For the sake of completeness, calculations carried out with the Dunning’s cc-pVXZ basis sets, $X=2,3,4,5$ being the cardinal number of the basis set, have been included in the Supplementary Information.
@@
Figure 2: Correlation energies ($E-E_{\mathrm{HF}}$) in mEh for the selected molecules obtained with GNOF (dashed), GNOFm (dotted), and CCSD(T) (solid) using $c$-pwCVXZ basis sets. The cardinal numbers X = 2, 3, and 4 correspond to blue, red, and green, respectively. Molecules ordered according to the numbering given in Fig. 1.
@@
1 Cyclopentadiene -1156.1 -1189.5 -1187.6
2 Pyrrole -1188.2 -1226.0 -1274.7
3 Furan -1221.9 -1259.2 -1302.5
4 Imidazole -1226.2 -1265.1 -1311.5
5 Benzene -1353.4 -1393.1 -1452.5
6 Pyridine -1388.5 -1432.1 -1489.6
7 Pyrazine -1425.6 -1471.6 -1528.4
8 Pyrimidine -1425.9 -1469.1 -1524.7
9 Pyridazine -1428.8 -1475.5 -1532.3
10 Triazine -1463.6 -1513.1 -1558.0
11 Thiophene -1542.9 -1587.3 -1579.2
12 Tetrazine -1506.4 -1554.9 -1612.3
@@
Table 1: Complete basis set (CBS) extrapolated correlation energies ($E-E_{\mathrm{HF}}$) in mEh for the 12 molecular systems, computed using GNOF, GNOFm, and CCSD(T). Helgaker’s extrapolation scheme, $E_{\infty}+bX^{-3}$, was employed with $X=2,3,4$ as the cardinal number of the basis set.
As shown in Fig. 2, energy differences for the same method and molecule decrease as the size of the basis sets is augmented. This suggests that we are approaching the CBS limit for the reported correlation energies. Therefore, we computed the CBS limit correlation energies for the twelve molecules set using GNOF and GNOFm, as well as CCSD(T). Previous studies [67] demonstrated similar results for different extrapolation schemes within the NOFT framework, concretely, exponential function like: $E(X)=E_{\infty}+Aexp(-\gamma X)$ and power function like: $E(X)=E^{\prime}_{\infty}+A^{\prime}X^{-\gamma^{\prime}}$, being $X$ the cardinal number of the basis set. Nevertheless, in this work we employ the Helgaker’s extrapolation scheme [61] of the form $E_{corr}=E^{{}^{\prime\prime}}_{\infty}+bX^{-3}$, which is considered a standard method today and is particularly adequate for Dunning correlation-consistent basis sets. The corresponding GNOF and GNOFm CBS estimated values are given in Table 3, which also presents extrapolations CCSD(T), performed using the same procedure. An inspection of CBS limit molecular correlation energies reveals an agreement within 100 mEh for GNOF in most cases, values that are even improved to around 50 mEh when GNOFm is utilized.
@@
Figure 3: Complete basis set (CBS) extrapolated correlation energies ($E-E_{\mathrm{HF}}$) in mEh for the 12 molecular systems, computed using GNOF, GNOFm, and CCSD(T). Helgaker’s extrapolation scheme, $E_{\infty}+bX^{-3}$, was employed with $X=2,3,4$ as the cardinal number of the basis set.
The results shown in Table 3 are summarized in Fig. S2. Here, we plot CBS extrapolated correlation energies for the 12 molecular systems. GNOFm energies systematically get closer to CCSD(T) results when the static term between electron pairs is modified according to Eq. (16). In other words, Fig. S2 reveals that GNOFm CBS correlation energies reduce differences between GNOF and CCSD(T) to the half.
@@
Both GNOF and its recent modification GNOFm provide qualitatively good descriptions of the five- and six-membered rings, as their traces lie close to and largely parallel with the CCSD(T) reference. Notably, Fig. S2 also shows a quantitative improvement of GNOFm over GNOF, consistent with Lew-Yee and Piris.[36] More importantly, the figure demonstrates that NOFs recover dynamic-correlation effects across the entire family of correlation-consistent Dunning basis sets and for all systems considered, proving the robustness of the Global NOF approach. Unlike earlier approximations, particularly among electron-pairing based NOFs described in section II, these functionals incorporate dynamic correlation within the energy expression itself and therefore do not require perturbative corrections.
Basis Set HF GNOF GNOFm
$c$-pwCVDZ 1.755 0.037 0.039
$c$-pwCVTZ 1.755 0.026 0.027
$c$-pwCVQZ 1.755 0.027 0.025
Table 2: Root mean square deviations (in a.u.) from experimental non-zero molecular dipole moment values computed using Hartree-Fock, GNOF and GNOFm, for Dunning’s $c$-pwCVDZ, cc-pwCVTZ and cc-pwCVQZ basis sets. Results for each molecule are given in the Supplementary Information, together with bibliographic references corresponding to experimental data.
Finally, in Table 4 we report root mean square deviations (in a.u.) from experimental data corresponding to non-zero molecular dipole moments. Results for each molecular system, method, and basis set, can be found in the Supplementary Information, as well as literature sources corresponding to experimental data. HF provides same errors for all studied basis sets. In fact, increasing the basis reduces representation error in HF, but it does not add correlation. Consequently, HF converges smoothly to the HF limit of dipole moments, which can differ from the correlated limit by a nearly basis-independent offset [68]. Previous studies [69] showed that dipole, quadrupole and octupole moments computed with electron-pairing-based NOFs compare with CCSD and MRSD-CI for small molecular systems. According to the results shown in Table 4, both GNOF and GNOFm improve significantly HF dipole moments, and their errors reduce with increasing the basis set, particularly for GNOFm.